### v2

# Quantifying Multilingual Performance of Large Language Models Across Languages

Abstract
The development of Large Language Models (LLMs) relies on extensive text corpora, which are often unevenly distributed across languages. This imbalance results in LLMs performing significantly better on high-resource languages like English, German, and French, while their capabilities in low-resource languages remain inadequate. Currently, there is a lack of quantitative methods to evaluate the performance of LLMs in these low-resource languages. To address this gap, we propose the Language Ranker, an intrinsic metric designed to benchmark and rank languages based on LLM performance using internal representations. By comparing the LLM’s internal representation of various languages against a baseline derived from English, we can assess the model’s multilingual capabilities in a robust and language-agnostic manner. Our analysis reveals that high-resource languages exhibit higher similarity scores with English, demonstrating superior performance, while low-resource languages show lower similarity scores, underscoring the effectiveness of our metric in assessing language-specific capabilities. Besides, the experiments show that there is a strong correlation between the LLM’s performance in different languages and the proportion of those languages in its pre-training corpus. These insights underscore the efficacy of the Language Ranker as a tool for evaluating LLM performance across different languages, particularly those with limited resources.
1 Introduction
Large Language Models (LLMs), such as GPT-4, Claude-3 and LlaMa-3, have demonstrated surprising performance in various NLP tasks Achiam et al. (2023); Ouyang et al. (2022); Touvron et al. (2023); Team et al. (2024); Jiang et al. (2023); Bai et al. (2023). However, the majority of the text datasets are presented in high-resource languages such as English Xie et al. (2024). According to the statistics, for GPT-3 model approximately 92.65% of the training tokens are English and all other languages share the remaining 7.35% training tokens OpenAI (2023). Similarly, English accounts for 89.70% of data for pre-training LlaMa 2 Touvron et al. (2023). This disparity leads to a bias where LLMs exhibit superior performance in high-resource languages like English, German, and French but struggle significantly with low-resource languages.
The imbalance in language representation during the training phase of LLMs means these models are less proficient in low-resource languages. For example, LLM cannot understand the true meaning of some slang terms with specific cultural background, such as Chinese idioms Zhang et al. (2023). Additionally, recent studies have shown that pre-trained models perform poorly in languages with insufficient training data Lankford et al. (2024). These observations highlight the need for a metric that quantitatively assesses LLM performance across different languages, especially for those languages with limited resources.
In this paper, we introduce Language Ranker, a novel method that leverages internal representations to quantitatively evaluate the multilingual capabilities of LLMs, particularly focusing on low-resource languages. We establish the representation of LLMs on the English corpus as a baseline and measure the similarity between this baseline and representations from other languages. This similarity metric serves as the performance score for each language. We validate our approach by applying the Language Ranker to five state-of-the-art open-source LLMs: LlaMa2, LlaMa3, Qwen, Mistral-v0.1, and Gemma. We also compare the Language Ranker’s performance with the proportion of each language in the training datasets and against other established benchmarks. Through comprehensive experiments, our major observations can be summarized as follows:
  • We demonstrate that high-resource languages show higher similarity scores with English, whereas low-resource languages exhibit lower similarity scores, validating the effectiveness of the Language Ranker in measuring language-specific performance.
  • We uncover a strong correlation between the LLM performance and the proportion of languages in the pre-training corpus, providing insights into the impact of training data distribution on LLM performance.
  • We show that high-resource languages are more evenly distributed in the embedding space, while low-resource languages tend to be narrowly clustered. This distribution analysis further supports our intrinsic metric’s reliability in assessing language performance.
2 The Proposed Method
In this section, we will give an introduction to our analysis method. First, we will introduce the dataset that we used in our experiment. Then, we will introduce how to obtain the similarity between English and other languages, as well as how to compare different LLMs’ performances.
2.1 Probing Datasets
We use OPUS-100 Zhang et al. (2020) as our evaluation datasets. OPUS-100 is an English-centric multilingual corpus that covers 100 languages. Each sample consists of text in a non-English language as the original data, with its English translation serving as the target data. For example, {"German": "Ich wollte dir erst noch etwas zeigen.","English": "I wanted to show you something first."}. After filtering, there are 94 subsets containing English, including high-resource languages such as German, French, and Chinese, as well as low-resource languages such as Oriya, Kannada, and Kazakh. Each subset contains 2000 samples.
2.2 Similarity Measurement
We employ cosine similarity to measure the LLMs’ performance gap between the target language and English. Specifically, given two sentences $X=\{x_{i}\}_{i=1}^{n}$ and $Y=\{y_{i}\}_{i=1}^{m}$ representing the text in English and the text in the target language. We use the representation obtained after LLM mapping of the last token $x_{n}$ and $y_{m}$ as the representation of the text and calculate the similarity between them. As we know, LLM consists of several layers of Transformer blocks Vaswani et al. (2017). Therefore, after each layer of mapping by the transformer block, we can get a representation vector $x_{n}^{l}$ and $y_{m}^{l}$, $l=1...H$, where $H$ represents the number of the layer of LLMs. According to Li et al. (2024), the intermediate representation can be briefly summarized by the following equations:
$x^{l+1}=\text{MLP}(x^{l}+\text{MHA}(x^{l}))\quad l=1...H,$ (1)
where MHA means multi-head attention or multi-group attention, and MLP means standard multilayer perceptron layer. Next, we take $x_{n}^{l}$ and $y_{m}^{l}$ to calculate the similarity. To implement a more robust similarity measure, we use the average similarity obtained by several intermediate layers as the final similarity. This process can be described as follows:
$Sim=\frac{1}{|l_{sub}|}\sum_{i=1}^{|l_{sub}|}Sim_{i},\,\text{where}\,Sim_{i}=\frac{x_{n}^{i}y_{m}^{i}}{||x_{n}^{i}||||y_{m}^{i}||},$ (2)
where $l_{sub}$ = $\{5,10,15,20,25\}$ is the subset of the layers we selected. Finally, we use $Sim$ to evaluate the performance gap between English and Non-English corpus.
2.3 Rank Correlation Measurement
[…]
Figure 1: Performance of different LLMs for ten kinds of language. High-resource languages: German, Spanish, French, Indonesian and Chinese; and low-resource languages: Igbo, Kazakh, Kannada, Oriya and Turkmen.
3 Experiments
In our experiments, we utilize five prominent open-source large models: LlaMa2 Touvron et al. (2023), LlaMa3 Meta-AI (2024), Qwen Bai et al. (2023), Mistral-v0.1 Jiang et al. (2023), and Gemma Team et al. (2024).
We conduct experiments to answer the following research questions: RQ1: Can the Language Ranker effectively quantify the performance of LLMs across multiple languages? (Section 3.1) RQ2: How consistent are the performance rankings of different LLMs when evaluated across a diverse set of languages? (Section 3.2) RQ3: Is the proposed cosine similarity metric correlated with the proportion of a language in the LLMs’ pre-training corpus? (Section 3.3) RQ4: Is the proposed cosine similarity metric correlated with performance on other benchmark tasks for quantifying the multilingual capabilities of LLMs? (Section 3.4)
3.1 Can Language Ranker Quantify LLM Performance Across Languages?
To visualize the performance of different LLMs in these languages, we selected 10 representative languages to display their inference results. They consist of five high-resource languages, including German, Spanish, French, Indonesian, and Chinese, and five low-resource languages, including Igbo, Kazakh, Kannada, Oriya, and Turkmen. Figure 1 shows detailed results, where the X-axis represents different layers of LLMs, while the Y-axis represents the similarity between the target language and English for each layer. From Figure 1, we can observe that high-resource languages have representations more similar to English, whereas low-resource languages show less similarity. Specifically, German, Spanish, French, and Malay generally maintain cosine similarity scores above 0.6, with Spanish and French often showing the highest scores, indicating that these languages are better represented in the models’ embeddings. In contrast, low-resource languages, such as Igbo, Kazakh, Kannada, Oriya, and Turkmen, display significantly lower cosine similarity scores, often below 0.4. These results show the disparities in performance across languages and highlights the utility of the Language Ranker in quantifying these differences robustly.
Figure 2: Rank correlation between different LLMs. This is calculated using metric introduced in Section 2.3. It shows high correlations across LLMs.
3.2 Comparison Across Different LLMs
In this section, we analyze how various LLMs perform across multiple languages using the Language Ranker. We focus on understanding the consistency of language performance among different LLMs, including models with varying architectures and training specifics. We have the following findings:
(1) Different models display similar results across languages. Figure 1 presents the cosine similarity scores across various layers for four different 7B parameter LLMs: LLaMa, Gemma, Mistral, and Qwen. Four LLMs display a similar trend where high-resource languages (i.e., German, Spanish, French, Malay, and Chinese) consistently exhibit higher cosine similarity scores compared to low-resource languages (i.e., Igbo, Kazakh, Kannada, Oriya, and Turkmen). Figure 2 further corroborates these findings by comparing the rank correlation of the similarity scores across the LLMs. Each LLM’s ranking is used as a baseline, and the remaining three models exhibit ranking patterns that are largely similar to this baseline. This similarity indicates that despite differences in model architecture or training specifics, the relative performance of languages remains consistent across these four models.
(2) Fine-tuning on specific languages will improve its performance. According to the technical report of Qwen Bai et al. (2023), Qwen has additional fine-tuning on the Chinese corpus, which leads to better performance in Chinese. In Figure 1, we observe that for LlaMa2, Gemma, and Mistral, the performance of Chinese is slightly lower than that of other high-resource languages. However, for Qwen, the performance of Chinese is roughly comparable to other high-resource languages and even shows a gradual improvement in the last few layers. This improvement is more clear in Qwen, mainly due to additional fine-tuning of the Qwen model family on the Chinese corpus, as noted in the technical report of Qwen.
Figure 3: Similarity scores of LlaMa2 7B and LlaMa3 8B. From left to right, the languages are Czech (cs), Korean (ko), German (de), Arabic (ar), Italian (it), Bengali (bn), French (fr), Turkish (tr), Turkmen (tk), Hindi (hi), Dutch (nl), Kazakh (kk), Kannada (kn), and Thai (th).
(3) Comparison of LlaMa2 and LlaMa3. We also explore the performance of LlaMa 2 7B and LlaMa 3 8B. From Figure 3, we can observe that in some high-resource languages, such as German, French, and Dutch, both LlaMa 2 and LlaMa 3 perform well, and LlaMa 3 performs better than LlaMa 2, for low-resource languages, both perform poorly, and LlaMa 3 performs worse than LlaMa 2. This phenomenon suggests that there is a certain consistency in the performance of high-resource languages and low-resource languages in similarity metrics, which can be used to distinguish high-resource and low-resource languages.
Language Proportion Similarity Language Proportion Similarity
German 0.17% 0.723 Welsh $\leq 0.01\%$ 0.396
French 0.16% 0.737 Persian $\leq 0.01\%$ 0.300
Swedish 0.15% 0.662 Urdu $\leq 0.01\%$ 0.275
Chinese 0.13% 0.552 Kannada $\leq 0.01\%$ 0.236
Table 1: The proportion of different languages in the LlaMa2 pre-training corpus and the similarity metric we proposed. The English language ratio is 89.7%.
Language ARC MMLU
LlaMa7B Gemma7B Mistral7B Qwen7B LlaMa7B Gemma7B Mistral7B Qwen7B
Chinese 27% 71% 57% 66% 32% 54% 37% 44%
German 27% 68% 63% 32% 25% 57% 47% 27%
French 31% 76% 59% 42% 24% 58% 48% 28%
Spanish 31% 77% 60% 46% 29% 56% 52% 33%
Italian 29% 77% 67% 44% 23% 56% 44% 32%
Kannada 24% 48% 27% 21% 21% 40% 22% 19%
Hindi 28% 60% 42% 22% 25% 45% 32% 23%
Armenian 19% 40% 36% 20% 20% 36% 30% 25%
Marathi 28% 46% 26% 25% 27% 42% 30% 26%
Telugu 30% 42% 30% 30% 24% 33% 34% 23%
Table 2: Performance on two inference tasks
Figure 4: Visualization of the embedding space of Gemma-7B for eight languages. Four figures at the top are high-resource languages, Four figures at the bottom are low-resource languages.
3.3 Relationship to Ratio of Training Corpus?
In this section, we explore the relationship between the proportion of each language in the LLaMa2 pre-training corpus and their corresponding performance as measured by the similarity metric. According to the technical report of LlaMa2 Touvron et al. (2023), we obtain the proportion of the pre-training corpus of some languages. Table 1 illustrates this relationship by listing a selection of languages with their proportion in the training data and their similarity scores relative to English. The table is divided into two parts: the left side lists high-resource languages with relatively higher proportions in the LLaMa2 pre-training corpus, and the right side lists low-resource languages with very low proportions ($\leq$ 0.01%). For example, German, with a proportion of 0.17%, has a high similarity score of 0.723, indicating strong performance in comparison to English. This trend suggests that languages with a higher proportion in the pre-training corpus tend to have higher similarity scores, reflecting better model performance. In contrast, low-resource languages like Kannada and Urdu, each with proportions of less than 0.01%, have much lower similarity scores (i.e., 0.236 and 0.275).
High-High Similarity Score High-Low Similarity Score Low-Low Similarity Score
English-German 0.72 German-Silesian 0.48 Azerbaijani-Turkmen 0.51
Italian-French 0.68 French-Erzya 0.35 Hungarian-Yiddish 0.24
German-French 0.67 Italian-Romany 0.32 Kab-SMT 0.36
French-Chinese 0.59 Italian-Uighur 0.27 Mari-Tatar 0.48
Table 3: Similarity score of different language pairs of Gemma-7b.
High-High Similarity Score High-Low Similarity Score Low-Low Similarity Score
English-German 0.72 German-Silesian 0.44 Azerbaijani-Turkmen 0.51
Italian-French 0.69 French-Erzya 0.31 Hungarian-Yiddish 0.19
German-French 0.68 Italian-Romany 0.15 Kab-SMT 0.40
French-Chinese 0.56 Italian-Uighur 0.20 Mari-Tatar 0.42
Table 4: Similarity score of different language pairs of llama-7b.
3.4 Correlation with Other Inference Tasks?
To more comprehensively reflect the performance of LLMs in various languages, we evaluate the multilingual reasoning ability of these LLMs.
We use MLMM-evaluation [^1] https://github.com/nlp-uoregon/mlmm-evaluation as our benchmark dataset to evaluate LLMs’ performances on reasoning tasks in various languages. The benchmark dataset can be used to evaluate the LLM across 26 different languages. It consists of three datasets: ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), and MMLU Hendrycks et al. (2020). We chose ARC and MMLU for evaluation, and both of them are multiple-choice datasets. The ARC dataset consists of 7,787 multiple-choice science questions drawn from a variety of sources. The MMLU dataset contains multiple-choice questions derived from diverse fields of knowledge. We selected five high-resource languages (Chinese, German, French, Spanish, Italian) and five low-resource languages (Kannada, Hindi, Armenian, Marathi, Telugu) for evaluation, randomly selected 100 samples from each language for 4-shot learning prediction, and used accuracy as the metric. The LLMs evaluated are consistent with Figure 1: LlaMa 7B, Gemma 7B, Mistral 7B, and Qwen 7B.
The predicted result is shown in Table 2. From the result, we find that for Gemma, Mistral, and Qwen, the performance of high-resource languages is significantly better than that of low-resource languages, and Gemma performs best. For the LlaMa, the performance in all languages is generally not as good as the first three LLMs. This result shows that LLM reasoning ability in low-resource languages is worse than that in high-resource languages. This result proves that there are differences in performance between high-resource and low-resource languages in reasoning tasks, illustrating the effectiveness of the proposed cosine similarity metric.
4 Further Analysis of Proposed Metric
In the last section, we introduced and evaluated the Language Ranker, demonstrating its ability to quantify the multilingual capabilities of LLMs by comparing their internal representations against an English baseline. This provided a robust measure of how LLMs perform across different languages, especially highlighting the disparities between high-resource and low-resource languages.
Building on these insights, in this section, we delve deeper into the proposed metric to explore its credibility and reliability further. Specifically, we aim to answer the following questions: RQ5: Is choosing English as the benchmark a wise choice (Section  4.1)? RQ6: What does the subspace of each language look like (Section  4.2)? RQ7: Is choosing cosine similarity a wise choice (Section  4.3)?
Figure 5: Box plot of the relationship between the double variance and similarity of Gemma 7B and Mistral 7B.
4.1 Why Using English as Baseline?
In the above sections, we choose English as a baseline. This is based on the a priori assumption that low-resource languages generally perform worse than high-resource languages. But if we choose other high-resource languages as baselines, will we get the same performance? In other words, how can we ensure that our metric is not affected by the English language itself? To answer this question, we divided our probing datasets into three types: High Resource-High Resource (H-H), High Resource-Low Resource (H-L), and Low Resource-Low Resource (L-L). To fulfill our requirement, we utilize Tatoeba-Challenge Tiedemann (2020) as our dataset instead of opus-100 because the latter is an English-centric dataset which means there is no Low Resource-Low Resource language pair. Tatoeba-Challenge is a challenge set for machine translation that contains 32G translation units in 2,539 bitexts. The whole data set covers 487 languages linked to each other in 4,024 language pairs. We select four language pairs for each group, English-German (en-de), English-French (en-fr), German-French (de-fr), and Italian-German (it-de) represent H-H; German - Silesian (de-szl), French-Erzya (fr-myv), Italian - Romany (it-ro) and Italian - Uighur (it-uig) represent H-L; Azerbaijani-Turkmen (az-tr), Hungarian-Yiddish (hu-yi), Kabyle-Standard Moroccan Tamazight (kab-SMT) and Mari-Tatar (ma-ta) represent L-L. The results are shown in Table 3 and Table 4.
From the results, we can observe that the score of High-High is higher than the score of High-Low and Low-Low universally. An obvious inference is that the distribution of high-resource languages is relatively close to each other, while the distribution of low-resource languages varies greatly, neither being close to each other nor to high-resource languages. Therefore, the distribution of high-resource languages is relatively consistent, while the distribution of low-resource languages varies greatly. Choosing English as the baseline is a convenient choice. Thus, we can also choose other high-resource languages such as German and French as the baseline.
4.2 Deeper Analysis of the Embedding Space
We explained why we chose English as the baseline in the above section. Choosing English is actually choosing a high-resource language as our baseline. A question naturally arises: How can we make sure that the performance of high-resource languages is better than the performance of low-resource languages? The result of Table 2 has confirmed the answer to the question by the reasoning task. To answer this question more deeply, we need to analyze the distribution of the embedding of different languages. As is shown in Figure 4, the top four sub-figures are embedding spaces of high-resource languages, while the bottom four sub-figures are embedding spaces of low-resource languages. It is obvious that high-resource languages are more evenly distributed throughout the space, while low-resource languages are more narrowly distributed, and compressed into a near straight line. Therefore, the performance of low-resource languages is worse than that of high-resource languages, which means it is suitable to choose a high-resource language like English as the baseline.
4.3 Why Using Cosine Similarity?
Recent research Steck et al. (2024) has shown that cosine similarity is not always a reliable metric. Inspired by Section 4.2, the quality of the performance of various languages can be clearly judged from the subspace distribution. Therefore, we decided to quantify the performance of these languages from a distribution perspective.
Back to Figure 4, the projection distributions of high-resource languages in different directions are relatively consistent, while the distribution of low-resource languages is compressed approximately into a straight line. The projection distributions outside the straight line, such as the projections perpendicular to the straight line, are crowded in a smaller area. This suggests that we can use the projection variance to approximately measure the quality of the distribution.
According to PCA, we assume the embedding vectors are $\{X_{i}\}_{i=1}^{n}$(being centralized), the projection direction is $\omega$, the project variance $Var(X,\omega)$ can be calculated as follows:
$Var(X,\omega)$ $=\frac{1}{n}\sum_{i=1}^{n}(X_{i}^{T}\omega)^{2}=\frac{1}{n}\sum_{i=1}^{n}\omega^{T}X_{i}X_{i}^{T}\omega$ (3)
$=\omega^{T}Cov(X)\omega\quad s.t.\quad\omega^{T}\omega=1$
It is obvious that $Var(X,\omega)$ is the eigenvalue of the $Cov(X)$. For high-resource languages, projection variance in different directions should be close to each other so that it can be evenly distributed in all directions. The opposite is true for low-resource languages. Therefore, we can extract the first K eigenvalues $\{Var(X,\omega_{i})\}_{i=1}^{k}$ and calculate their variance. The variance of the eigenvalues can be used to measure the differences in the distribution in each direction, which is called double variance. This metric can be used to specifically measure the quality of the distribution. The higher the double variance, the more unbalanced the distribution and the worse the performance, and the vice versa.
Gemma-7b Mistral-7b
High-resource Low-resource High-resource Low-resource
Language Double Var Similarity Language Double Var Similarity Language Double Var Similarity Language Double Var Similarity
Italian 0.04 0.66 Nepali 0.75 0.42 Italian 0.12 0.59 Nepali 0.60 0.28
French 0.09 0.69 Kazakh 0.85 0.38 French 0.17 0.65 Kazakh 0.32 0.32
Spanish 0.06 0.68 Burmese 0.36 0.30 Spanish 0.17 0.64 Burmese 0.40 0.20
German 0.10 0.72 Pashto 0.72 0.36 German 0.19 0.66 Pashto 0.73 0.29
Table 5: Similarity scores and double variance results for some languages on Gemma-7b and Mistral-7b.
We employ the box plot to show the relationship between the proposed cosine similarity metric and the double variance metric more clearly. From the results of Table 5, we can observe that for each LLM, the languages in the left part are some common high-resource languages, which have higher similarity and lower double variance, while the right part is the opposite for low-resource languages. The second observation is that as the variance increases, the similarity score also tends to decrease. The increase in variance means that the distribution of the subspace becomes uneven and the similarity score decreases accordingly. This shows that the proposed cosine similarity metric can be utilized to roughly measure the quality of distribution of the subspace, which can thus measure the performance of LLM in different languages.
5 Related Work
In this section, we summarize two lines of research that are most relevant to ours.
Representation Engineering.  Representation engineering has emerged as an effective approach to enhance the interpretability and transparency of LLMs. Researchers have been leveraging internal representations to tackle various challenges. Zou et al. (2023) summarizes the application of representation engineering in bias, fairness, model editing, and other areas. Schwenk (2007) found that the internal representation of LLM has a certain correlation with time and space, and the internal representation can be employed to represent time and space. Li et al. (2024) found that the representation of the attention head inside the LLM can be used to indicate the correct reasoning direction, probe analysis is further used to correct the internal representation direction to improve the LLM’s performance. Marks and Tegmark (2023) study the structure of LLM representations of true/false statements, proved that language models linearly represent the true/false of factual statements. Ju et al. (2024) used probe technique to detect how LLM stores knowledge layer by layer.
Multilingual Language Model.  Recent research such as Qin et al. (2024) summarizes the recent progress and future trends in multilingual large language models. Ahuja et al. (2024) constructed a benchmark to evaluate LLM’s multilingual ability comprising 22 datasets covering 83 languages. Huang et al. (2023); Qin et al. (2023) have proven that LLM performance varies substantially across different languages and they employ a prompt technique to improve task performance across languages. Wendler et al. (2024) explores how LlaMa works in multilingual tasks and what role English plays in these tasks. The imbalance distribution of training corpus in different languages leads to the bias of LLM towards some high-resource languages such as English Blasi et al. (2021). Some approaches employ multilingual language modeling to alleviate the phenomenon Shen et al. (2024); Kalyan et al. (2021); Conneau et al. (2019). These studies show the importance of strengthening the cross-lingual capabilities of the pre-trained model. Schäfer et al. (2024) found that the presence of a primary language in the training process of LLMs can improve the performance of low-resource languages and lead to a more consistent representation of LLMs in different languages. Liu et al. (2024) found that for English-centric LLMs, although translation into English helps improve the performance of NLP tasks, it is not the best choice for all situations.
6 Conclusions and Future Work
In this work, we propose the Language Ranker to evaluate the performance of LLMs across diverse languages by comparing their internal representations to English. The results show that high-resource languages show higher similarity scores with English, while low-resource languages have lower scores, validating the effectiveness of our metric in assessing language performance. Besides, there is a strong correlation between the performance of LLMs in different languages and the proportion of those languages in the pre-training corpus. Further, results indicate that high-resource languages are more evenly distributed in the embedding space, whereas low-resource languages tend to be narrowly clustered. In the future, we plan to design more comprehensive benchmarks to measure LLM’s capabilities in different languages.
Limitations
The proposed Language Ranker approach provides an initial quantitative way to analyze LLM performance across languages. We acknowledge that the language ranker method we proposed offers only a rough measurement. While our findings indicate a correlation between the similarity scores and the proportion of each language in the pre-training dataset, these scores alone are not sufficient to precisely measure the exact proportions. Our intent was to provide an initial quantitative approach to explore this relationship, and we recognize the need for more comprehensive methods and additional metrics to accurately assess the impact of pre-training data distribution across languages. Furthermore, the method does not account for potential biases or skews that could be present in the multilingual evaluation datasets themselves. The existence of such biases can also introduce noise in the resulting rankings of language abilities for different LLMs.
Appendix A Appendix
A.1 Ranking Result For LLMs
We give the similarity scores of the four LLMs used in the experiment on 18 high-resource languages. Results are shown in the following tables.
Language Similarity Score Language Similarity Score
German 0.723 Western Frisian 0.378
French 0.737 Tamil 0.347
Spanish 0.768 Gujarati 0.313
Italian 0.706 Kurdish 0.308
Russian 0.734 Pashto 0.284
Dutch 0.709 Assamese 0.260
Polish 0.664 Central Khmer 0.240
Malay 0.651 Panjabi 0.218
Swedish 0.661 Amharic 0.202
Table 6: The similarity score of LlaMa2 7B.
Language Similarity Score Language Similarity Score
German 0.719 Western Frisian 0.443
French 0.691 Tamil 0.420
Spanish 0.683 Gujarati 0.433
Italian 0.662 Kurdish 0.358
Russian 0.674 Pashto 0.362
Dutch 0.658 Assamese 0.396
Polish 0.618 Central Khmer 0.330
Malay 0.615 Panjabi 0.379
Swedish 0.629 Amharic 0.298
Table 7: The similarity score of Gemma 7B.
Language Similarity Score Language Similarity Score
German 0.639 Western Frisian 0.346
French 0.623 Tamil 0.279
Spanish 0.616 Gujarati 0.270
Italian 0.571 Kurdish 0.262
Russian 0.611 Pashto 0.267
Dutch 0.566 Assamese 0.276
Polish 0.514 Central Khmer 0.252
Malay 0.497 Panjabi 0.213
Swedish 0.532 Amharic 0.191
Table 8: The similarity score of Mistral 7B.
Language Similarity Score Language Similarity Score
German 0.805 Western Frisian 0.441
French 0.793 Tamil 0.510
Spanish 0.800 Gujarati 0.469
Italian 0.773 Kurdish 0.436
Russian 0.794 Pashto 0.448
Dutch 0.773 Assamese 0.507
Polish 0.752 Central Khmer 0.407
Malay 0.730 Panjabi 0.385
Swedish 0.759 Amharic 0.470
Table 9: The similarity score of Qwen 7B.
Figure 6: Simlarity scores curves of LlaMa-13b and Qwen-13b.
Figure 7: Simlarity scores curves of Bloom-3b and Bloom-7b.
A.2 Details of experiment in Table 4 and Table 3
We selected data from the Tatoeba-Challenge repository [^2] https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/data. Since the number of samples for some low-resource language pairs is small, we extracted 100 samples for each language pair. If there are less than 100, we extracted all samples and extracted them according to the test-dev-train priority.
High-High Similarity Score High-Low Similarity Score Low-Low Similarity Score
English-German 0.64 German-Silesian 0.43 Azerbaijani-Turkmen 0.51
Italian-French 0.60 French-Erzya 0.30 Hungarian-Yiddish 0.18
German-French 0.62 Italian-Romany 0.25 Kab-SMT 0.39
French-Chinese 0.57 Italian-Uighur 0.24 Mari-Tatar 0.46
Table 10: Similarity score of different language pairs of mistral-7b.
High-High Similarity Score High-Low Similarity Score Low-Low Similarity Score
English-German 0.81 German-Silesian 0.58 Azerbaijani-Turkmen 0.67
Italian-French 0.75 French-Erzya 0.51 Hungarian-Yiddish 0.39
German-French 0.80 Italian-Romany 0.53 Kab-SMT 0.55
French-Chinese 0.71 Italian-Uighur 0.49 Mari-Tatar 0.55
Table 11: Similarity score of different language pairs of qwen-7b.
A.3 Performance of LLMs with Other Sizes
A.3.1 LlaMa-13b and Qwen-13b
We also explore the performance of the similarity metric in LLM of 13B parameters, We use LlaMa-13b and Qwen-13b to display the results. From Figure 6 we can observe that the partial order results of LLM-13B are roughly the same as those of LLM-7B, and there is a clear gap between high-resource languages and low-resource languages.
A.3.2 Bloom-3b and Bloom-7b
Figure 7 shows the result of Bloom-3b and Bloom-7b. Except for the last few layers of the model, there are obvious differences between high-resource languages and low-resource languages which are similar to the above LLMs, while there are smaller differences within the same category of languages. The score is higher than LlaMa, Gemma, Mistral, and Qwen.